7m+12n=-1
5m-3n=7

Answer by mananth(16946)

7 m + 12 n = -1 .............1
Total value
5 m -3 n = 7 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 4
7 m 12 n = -1
20 m -12 n = 28
Add the two equations
27 m = 27
/ 27
m = 1
plug value of m in (1)
7 m + 12 n = -1
7 + 12 n = -1
12 n = -1 -7
12 n = -8
n = - 2/3